OpenStudy (anonymous):

What is the y-intercept of the graph of -2=x-4? Help

OpenStudy (anonymous):

-4

OpenStudy (tamtoan):

to find y-intercept, let x=0 and solve for y

OpenStudy (anonymous):

there is no y to solve for

OpenStudy (anonymous):

X is 1

OpenStudy (anonymous):

it's a slope intercept equation y=mx+b

OpenStudy (anonymous):

because it is by it's self

OpenStudy (anonymous):

m=1 b=-4 m=slope b=y-intercept

OpenStudy (anonymous):

Hold on let me try it and see what i get Im studying for the match ECA's tomorrow

OpenStudy (anonymous):

Im not sure why there is a -2 in the 'y' spot though...but there is no y to solve for

OpenStudy (anonymous):

How did you get 1 what about negative 2 what did you do to that

OpenStudy (anonymous):

at first glance the problem you wrote looks like a slope intercept equation which is y=mx+b with x and y being a point (x,y) m is the variable for slope and b is the variable for y-intercept

OpenStudy (anonymous):

so it looked like your slope was 1 (because of the understood one in front of an x) and your y-intercept was -4

OpenStudy (anonymous):

But I do not know what to make of the -2 where the 'y' would be. Are you sure you typed it out correctly

OpenStudy (anonymous):

yes i did it looks like this -2y=x-4

OpenStudy (anonymous):

OH there is the missing y! divide all terms by -2

OpenStudy (anonymous):